OpenStudy (anonymous):

which is the exact value of the expression 6 sqrt(x)-sqrt(4v^2)-sqrt(36v)+sqrt(v^2)

OpenStudy (anonymous):

$$\sqrt{4v^2}=2v,\sqrt{v^2}=v,\sqrt{36v}=6\sqrt{v}$$

OpenStudy (anonymous):

So we have \(6\sqrt{x}-2v-6\sqrt{v}+v=-v-6\sqrt{v}+6\sqrt{x}\)

OpenStudy (anonymous):

this are my possible answers 3v \[v \sqrt{v}\] -v \[6v ^{2}\]

OpenStudy (anonymous):

@msellier that's not meant to be an \(x\), then, is it? :-)

OpenStudy (anonymous):

If you meant \(6\sqrt{v}-\sqrt{4v^2}-\sqrt{36v}+\sqrt{v^2}=6\sqrt{v}-2v-6\sqrt{v}+v=-v\)

OpenStudy (anonymous):

yes sorry my bad i meant a v thx for your help now i got it thank you

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

could you help me with another on please